X-84=4*(X-84)
With the above equations How do i know the value of X.Thanks

Answer by stanbon(75887)

X-84=4*(X-84)
Use the distributive law on the right side to get:
x-84 = 4x-336
Subtract x from both sides; Add 336 to both sides:
3x=252
Divide by 3 to get
x=84
------------
Cheers,
Stan H.